An urn contains 7 white balls and 3 black balls. Two balls are selected at random without replacement. What is the probability :1 The first ball is black and the second ball is white.?
2: One ball is white and the other is black?
3:the two balls are white ?
Answer #1

An urn contains

  • 7 white balls and
  • 3 black balls

The total number of balls is therefore 10

We draw two balls without replacement

1.

P(the first ball is black and the second ball is white)

= P(the first ball is black) * P(the second ball is white/first ball is black)

= (3/10) * (7/9)

= 7/30

= 0.2333

2.

P(One ball is white and the other is black)

= P(the first ball is black) * P(the second ball is white/first ball is black) + P(the first ball is white ) * P(the second ball is black/first ball is white)

= (3/10 * 7/9)+(7/10*3/9)

=7/15

=0.4667

3.

P(Two balls are white)

= P(the first ball is white) * P(the second ball is white/first ball is white)

= 7/10 * 6/9

=7/15

=0.4667

Let me know in the comments if anything is not clear. I will reply ASAP! Please do upvote if satisfied!